OpenStudy (anonymous):

What is the first step in solving the logarithmic equation, log 3 + log (x+1) = log (2x+4)?

Nnesha (nnesha):

quotient rule\[\huge\rm log_b y - \log_b x = \log_b \frac{ x }{ y}\] to condense you can change subtraction to division product rule \[\huge\rm log_b x + \log_b y = \log_b( x \times y )\] addition ----> multiplication power rule \[\huge\rm log_b x^y = y \log_b x\]

Nnesha (nnesha):

first step) condense look at the log properties there is "+" sign how would you write it in single log form ?

OpenStudy (xapproachesinfinity):

if \[\log a=\log b \Longrightarrow a=b\] after using what @Nnesha mentioned you can use this to get the solution